-6(x+8)+7(x+5)= 8x-5+6x-8
Answer by robertb(5830)

-6x - 48 + 7x + 35 = 14x - 13
x - 13 = 14x - 13
-13 + 13 = 13x
0 = 13x
x = 0/13 = 0.